# Online Navigation Refinement: Achieving Lane-Level Guidance by Associating Standard-Definition and Online Perception Maps

### v3

@@ B.1 Motivation
What is the relationship between this paper and planning methods? The task and methodology of this paper focus on delivering road-level navigation on vehicles and drivers’ devices. Addresses the issue of road-level navigation provision in a context where HD maps are no longer used. Recently, research like NavigScene peng2025navigscene has incorporated SD navigation data into autonomous driving planning modules. We suggest that our approach can act as a preliminary stage for NavigScene, enhancing the SD navigation data to offer lane-level guidance. This grants NavigScene’s planning module more detailed navigation details, thereby increasing planning precision.
@@ C.1 End-to-end Autonomous Driving
Planning, being an essential component for numerous fields zhang2025investigating; zhang2024exploring; zhang2025fully; bai2024learning, is recognized as one of the pivotal research domains in the realm of autonomous driving at present. VAD jiang2023vad utilizes an ego query mechanism to forecast individual mode trajectories, whereas VADv2 chen2024vadv2 enhances this by adopting a probabilistic framework that considers multiple trajectories. SparseDrive sun2025sparsedrive is innovative in its creation of parallel motion and planning modules, effectively decreasing the computational requirements of the BEV features. DiffusionDrive liao2025diffusiondrive employs a truncated diffusion policy to improve the probabilistic depiction of trajectories, while MomAD song2025don aims to enhance stability and maintain consistency throughout sequential planning decisions. Significantly, unlike other end-to-end planning techniques, NavigScene peng2025navigscene uses navigation data based on SD maps, achieving more accurate and reliable autonomous driving, thus underscoring the significance of navigation data. Consequently, we propose that the refinement of lane-level navigation through MAT could further enhance the capabilities of current autonomous driving methodologies.
@@ D.2 Range of Sample
The model is trained on nuScenes using synchronized LiDAR and camera inputs with official configuration. For each sample, an SD map cropping measure $150\,m\times 150\,m$ centered around the ego vehicle preserves the adjacent topological context followed by the sensor setting of nuScenes caesar2020nuscenes. For the OP map, the cropping measure of the OP map $30\,m\times 60\,m$ was centered around the vehicle of the ego, as referenced in maptrv2; liu2023vectormapnet; li2022hdmapnet.
@@ D.3 Pon split and OP maps
Both Train, val and test set apply the pon split roddick2020predicting of the nuScenes dataset nuscenes2019, ensuring that there is no leakage between the training and validation datasets. For consistent lane prediction segmentation with OMA test set, we re-trained MapTRv2 maptrv2 using the nuScenes dataset with the pon split. Drawing inspiration from the private protocol in MOT17/MOT20 dendorfer2021motchallenge, we suggest that future research evaluates the test set with an enhanced centerline prediction network, without relying on MapTRv2 as a baseline. Furthermore, we will establish an open evaluation protocol allowing submissions to include their own OP maps, thereby reducing reliance on a single model.
@@ E.1 Details of NR-PR
where thth are the thresholds in association P-R as $[0.5:0.05:0.95]$ (10 thresholds). Additionally, in the validation set, because NR-R is always 1, we use the NR-P value directly as NR-F1.
@@ G.1 Hyper Parameter
Length intervals of NR P-R. To further investigate the influence of the path length distribution on the NR-F1 metric, we performed a stratified evaluation based on length intervals. The motivation for introducing length intervals stems from our observation that within the dataset, short-distance paths occur far more frequently than long-distance ones, and their corresponding F1 scores tend to be higher. Without appropriate handling, this imbalance could bias the overall NR-F1 score toward performance on short-distance paths. Inspired by the size‑based categorization strategy adopted in the MS COCO benchmark lin2014microsoft, we categorize paths into discrete length intervals, compute the F1 score within each interval, and subsequently aggregate the results into a unified NR-F1 measure. This procedure ensures that each length range contributes proportionally to the final metric, thus mitigating the dominance of short‑distance paths in the evaluation. As shown in Tab. 18, omitting this stratification (that is, computing the metric on all paths without length‑based grouping) produces a substantially inflated overall score. In contrast, increasing the number of length intervals causes the aggregated metric to gradually decrease and eventually converge to a stable value. These findings suggest that the length interval-based computation allows NR-F1 to provide a more balanced and fair evaluation on both the short- and long-distance paths, resulting in a more comprehensive evaluation of the model’s ability to capture relational associations.
@@ G.3 Robustness under Environmental Noise and Occlusion
Robustness against Environmental Noise. We utilized nuScenes-C xie2025benchmarking, a benchmark designed to assess robustness against input corruptions, to simulate sensor degradation. Specifically, we used MapTRv2 maptrv2 to perform inference on the nuScenes-C validation set with pon split, generating OP maps under three specific weather conditions: Rain, Fog and Dark (Night). These conditions were categorized into three difficulty levels: Easy, medium, and Hard followed by the nuScenes-C setting.
@@ H.4 Generalization on Different OP Generators
To further validate the robustness of MAT against varying noise patterns inherent to different map generation paradigms, we visualize the association results on the OMA Test set using three distinct baselines: MapTR (liao2023maptr), MapTRv2 (maptrv2), and SeqGrowGraph (SSG) (xie2025seqgrowgraph).

### v4

@@ 7 Conclusion
Acknowledgments
This work was supported by the National Natural Science Foundation of China (62433003, 62476017).
@@ B.1 Motivation
What is the relationship between this paper and planning methods? The task and methodology of this paper focus on delivering road-level navigation on vehicles and drivers’ devices. Addresses the issue of road-level navigation provision in a context where HD maps are no longer used. Recently, research like NavigScene Peng et al. (2025) has incorporated SD navigation data into autonomous driving planning modules. We suggest that our approach can act as a preliminary stage for NavigScene, enhancing the SD navigation data to offer lane-level guidance. This grants NavigScene’s planning module more detailed navigation details, thereby increasing planning precision.
@@ C.1 End-to-end Autonomous Driving
Planning, being an essential component for numerous fields Zhang et al. (2025a; 2024c; b); Bai et al. (2024), is recognized as one of the pivotal research domains in the realm of autonomous driving at present. VAD Jiang et al. (2023) utilizes an ego query mechanism to forecast individual mode trajectories, whereas VADv2 Chen et al. (2024) enhances this by adopting a probabilistic framework that considers multiple trajectories. SparseDrive Sun et al. (2025) is innovative in its creation of parallel motion and planning modules, effectively decreasing the computational requirements of the BEV features. DiffusionDrive Liao et al. (2025) employs a truncated diffusion policy to improve the probabilistic depiction of trajectories, while MomAD Song et al. (2025) aims to enhance stability and maintain consistency throughout sequential planning decisions. Significantly, unlike other end-to-end planning techniques, NavigScene Peng et al. (2025) uses navigation data based on SD maps, achieving more accurate and reliable autonomous driving, thus underscoring the significance of navigation data. Consequently, we propose that the refinement of lane-level navigation through MAT could further enhance the capabilities of current autonomous driving methodologies.
@@ D.2 Range of Sample
The model is trained on nuScenes using synchronized LiDAR and camera inputs with official configuration. For each sample, an SD map cropping measure $150\,m\times 150\,m$ centered around the ego vehicle preserves the adjacent topological context followed by the sensor setting of nuScenes Caesar et al. (2020a). For the OP map, the cropping measure of the OP map $30\,m\times 60\,m$ was centered around the vehicle of the ego, as referenced in Liao et al. (2023b); Liu et al. (2023a); Li et al. (2022).
@@ D.3 Pon split and OP maps
Both Train, val and test set apply the pon split Roddick and Cipolla (2020) of the nuScenes dataset Caesar et al. (2020b), ensuring that there is no leakage between the training and validation datasets. For consistent lane prediction segmentation with OMA test set, we re-trained MapTRv2 Liao et al. (2023b) using the nuScenes dataset with the pon split. Drawing inspiration from the private protocol in MOT17/MOT20 Dendorfer et al. (2021), we suggest that future research evaluates the test set with an enhanced centerline prediction network, without relying on MapTRv2 as a baseline. Furthermore, we will establish an open evaluation protocol allowing submissions to include their own OP maps, thereby reducing reliance on a single model.
@@ E.1 Details of NR-PR
where $th$ are the thresholds in association P-R as $[0.5:0.05:0.95]$ (10 thresholds). Additionally, in the validation set, because NR-R is always 1, we use the NR-P value directly as NR-F1.
@@ G.1 Hyper Parameter
Length intervals of NR P-R. To further investigate the influence of the path length distribution on the NR-F1 metric, we performed a stratified evaluation based on length intervals. The motivation for introducing length intervals stems from our observation that within the dataset, short-distance paths occur far more frequently than long-distance ones, and their corresponding F1 scores tend to be higher. Without appropriate handling, this imbalance could bias the overall NR-F1 score toward performance on short-distance paths. Inspired by the size‑based categorization strategy adopted in the MS COCO benchmark Lin et al. (2014), we categorize paths into discrete length intervals, compute the F1 score within each interval, and subsequently aggregate the results into a unified NR-F1 measure. This procedure ensures that each length range contributes proportionally to the final metric, thus mitigating the dominance of short‑distance paths in the evaluation. As shown in Tab. 18, omitting this stratification (that is, computing the metric on all paths without length‑based grouping) produces a substantially inflated overall score. In contrast, increasing the number of length intervals causes the aggregated metric to gradually decrease and eventually converge to a stable value. These findings suggest that the length interval-based computation allows NR-F1 to provide a more balanced and fair evaluation on both the short- and long-distance paths, resulting in a more comprehensive evaluation of the model’s ability to capture relational associations.
@@ G.3 Robustness under Environmental Noise and Occlusion
Robustness against Environmental Noise. We utilized nuScenes-C Xie et al. (2025b), a benchmark designed to assess robustness against input corruptions, to simulate sensor degradation. Specifically, we used MapTRv2 Liao et al. (2023b) to perform inference on the nuScenes-C validation set with pon split, generating OP maps under three specific weather conditions: Rain, Fog and Dark (Night). These conditions were categorized into three difficulty levels: Easy, medium, and Hard followed by the nuScenes-C setting.
@@ H.4 Generalization on Different OP Generators
To further validate the robustness of MAT against varying noise patterns inherent to different map generation paradigms, we visualize the association results on the OMA Test set using three distinct baselines: MapTR (Liao et al., 2023a), MapTRv2 (Liao et al., 2023b), and SeqGrowGraph (SSG) (Xie et al., 2025a).